Inc. uses a perpetual inventory system. The following is information about one inventory item for the month of September:

Sep. 1

Inventory

20 units at $20

Sep. 4

Sold

10 units

Sep. 10

Purchased

30 units at $25

Sep. 17

Sold

20 units

Sep. 30

Purchased

15 units at $30

My answers in red*

If inc. uses the FIFO inventory method, what will be the remaining unit cost and total dollar amount of the 35 units still in Inventory on the Balance Sheet as of September 30?

Remaining unit cost:

__20___units at _$25___     and ____15___Units at $30_      

Total dollar amount: $950

If inc. uses the LIFO inventory method, what will be the remaining unit cost and total dollar amount of the 35 units in Inventory on the Balance Sheet?

Remaining unit cost:

____10__Units at $20   and 10 units at $25 and 15 units at $30

Total dollar amount: $900

My answers are bold, not sure if I approached this right! please help.
Answer #1

Total Units available for sale = 20+30+15 = 65 Units

Sale unit = 10+20 = 30 Units

a) FIFO ending inventory = 20 Units at $25 and 15 Units at $30 = 950

b) LIFO ending inventory = 10 Units at $20; 10 Units at $25 and 15 Units at $30 = $900